if 17 is added to the number im thinking of and then the sum is multiplied by 25, i get 750. what number am i thinking

Answer by Adebimpe(3)

Let the number am thinking of be x; x
i add 17 to the number; x + 17
i multiplied the result by 25; (x +17)25
the result is 750; (x + 17)25 = 750
25x + 425=750
25x =750-425
25x = 325
x=13